EVA Report – March 6th

EVA Report:

EVA #12 Crew members: Alexandre Martin (EVA leader), Gabriel Payen, Jérémy AUCLAIR, Victoria Da Poian, Axel Monnier (journalist), Bertrand Guez (journalist)

Author: Alexandre Martin


Near Pooh’s Corner: 12S-518900E-4250900N

Lith Canyon: 12S-518300E-4256500N

Time: departure at 9:30 a.m. Duration: 3h


  • Replace the LOAC
  • Explore Lith Canyon


We made some shots with the journalist in front of the Hab, and then we left the Hab to replace the LOAC at Pooh’s corner. We went to Lith Canyon to show our journalists a Martian-type environment. They took some shots with their drone, and we came back at the Hab, where they took some final shots.

9:30 am: EVA#12 left airlock

9:45 am: EVA#12 left the Hab

9:46 am: EVA#12 arrived at Pooh’s corner

10:05 am: EVA#12 left Pooh’s Corner

10:30 am: EVA#12 arrived at Lith Canyon

11:50 am: EVA#12 left Lith Canyon

12:15 pm: EVA#12 back at the Hab

12:30 pm: EVA#12 back in the Hab

EVA Report – March 5th


Here is today’s EVA report also comming from Jérémy:

EVA Report:
EVA #11 Crew members: Jérémy Auclair (EVA leader), Louis Mangin, Gabriel Payen, Benoit Floquet

– Around the Hab
– Near Pooh’s Corner: 12S-518900E-4252000N
– Near Mid Ridge Planitia: 12S-517500E-4251500N

Time: departure at 09:12 a.m. Duration: 2 hour 24 minutes

When describing shortly today’s EVA, we could say that concerning our experiments it was a failure. After arriving at the instruments’ position near Pooh’s corner, we found out that both batteries were empty. Therefore, we plugged the new battery on MegaARES to start transferring data on Gabriel’s USB key. We took one empty battery (to start charging it) and went back to the Hab to start Benoit’s localization experiment. After more than five minutes the GPS chip still hadn’t found any signal.

We decided to continue the walk to get on the Hab ridge. We took quite a few pictures, Louis with his camera and me with mine, on which I had mounted my orange filter (for my personal pictures). After we arrived over the Hab we headed back to the Hab. We took the ATVs to go back to Pooh’s Corner to finish MegaARES’ data transfer.

Once the transfer was done we plugged the LOAC on the battery. But once again the power system was broken. This time it was the switch which was broken. We took the decision to bring the two batteries and the LOAC back to the Hab to fix the power system.

Unfortunately, on the way back to the Hab, one of the batteries fell of one of the ATV 30 meters away of the Hab because of a bump on the road. It broke and started to leak slightly. We must have attached it too weakly because we had to bring back two batteries instead of one. I quickly asked Habcom to put a garbage bag in the engineering Airlock. 3 Minutes later we put the leaking battery in the plastic bag to confine it. Once this (small) crisis was taken care of we started the final engineering check and ended the EVA.

– LOAC: Test of the battery level
– MegaARES: Replacement of the battery
– Localization experiment: test with increasing uncertainty on the positioning measure

Vehicles: 4 ATVs

9:12 a.m.: EVA#11 left Airlock
9:20 a.m.: EVA left the Hab using the four ATVs
9:33 a.m.: Arrival on the working place and beginning of battery tests and data transfert
9:39 a.m.: Return to the ATVs
9:42 a.m.: Turning on the ATVs to go back to the Hab
9:46 a.m.: Arrival to the Hab, Benoit starts his GPS instrument 9:50 a.m.: Beginning of Walk to find a way to climb on the ridge 9:57 a.m.: Benoit’s instrument still doesn’t work
10:22 a.m.: Louis finds a safe way to climb on the ridge
10:32 a.m.: Arrival on top of the ridge
10:42 a.m.: Beginning of way down the ridge
11:00 a.m.: Departure to the instruments on Pooh’s Corner
11:18 a.m.: Departure to the Hab with both batteries and the LOAC 11:36 a.m.: Entry in the Main Airlock



EVA Report – March 1st

EVA Report:
EVA #8 Crew members: Alexandre Martin (EVA leader), Benoît FLOQUET, Laurent BIZIEN, Jérémy AUCLAIR

Author: Alexandre Martin

Near Pooh’s Corner: 12S-518900E-4250900N
Lith Canyon: 12S-518300E-4256500N

Time: departure at 9:00 p.m. Duration: 3h

• Replace the battery of the LOAC
• Explore Lith Canyon

We started our EVA placing Laurent’s solar panels experiment next to the GreenHab. After a small trip to Pooh’s Corner to deal with our instruments, we rode our ATV to explore Lith Canyon. We made a large 1h40 walk there, enough to visit a great part of it. We then went back to the Hab, yet we were not able to contact our Habcom. We discovered later that this was a walkie-talkie issue, maybe because of the weather (fast winds). 10 minutes later, he was finally able to hear us, and we could come back in the Hab

9:00 am: EVA#8 left airlock
9:15 am: EVA#8 left the Hab
9:16 am: EVA#8 arrived at Pooh’s corner
9:25 am: EVA#8 left Pooh’s Corner
9:45 am: EVA#8 arrived at Lith Canyon
11:25 am: EVA#8 left Lith Canyon
11:45 am: EVA#8 back at the Hab
12:00 am: EVA#8 back in the Hab

EVA Report – February 27th

EVA Report:

EVA #7 Crew members: Laurent BIZIEN (EVA leader), Benoît FLOQUET, Louis MANGIN, Gabriel PAYEN, Victoria DA-POIAN

Author: Laurent BIZIEN


Near Pooh’s Corner: 12S-518900E-4250900N

Hab surroundings

Time: departure at 2:12 p.m. Duration: 1h30 (shortened by 1h15)


  • Replace the battery of MegaARES
  • Train to emergency procedures


We started the EVA by splitting the involved members of the crew in two teams (we have learnt later by Shannon that it wasn’t something to do in EVA so I take the responsibility of such a decision error). Team red, composed of Victoria, Gabriel and Benoît took a Rover, Curiosity, and an ATV, 350.2. They went to Pooh’s Corner in order to change the battery of MegaARES and also checked the battery of LOAC instrument. At the same time, team blue, Louis and I, did the usual checks of the other Rovers (today, only Opportunity was there) and ATVs and of the water tanks. We also tried to inflate the tires of ATV 350.1. We didn’t succeed for the front left tire.

Team red quickly returned and we could start the emergency procedures. On the morning, I had made a speech to the other members of the crew about the emergency procedures some future marsonauts could encounter in EVA.

We started with a basic situation. One person simulating a health problem and another person to make a review. After a few instances, we followed up on emergency transportation. To do so, we brought Curiosity next to a person simulating a victim. The exercise’s goal was to transport the victim up to the Rover, from different ways depending of the victim’s injury. Finally, we performed realistic injury cases (bone traumatism, unconsciousness, respiratory arrest…) leading to emergency transportation and LSPs in the Rover’s trunk.

Realizing that the emergency procedures in EVA were very limited due to the lack of ability to move of the backpacks and of the helmets, we decided to return to the Hab. We carried Victoria into the main airlock and Gabriel stayed with her while Louis, Benoît and I went back to the Hab by the engineering airlock in order to simulate an emergency return.

2:12 p.m.: EVA#7 left Airlock

2:15 p.m.: Team red left the Hab

2:20 p.m.: Team red’s arrival at Pooh’s corner

2:30 p.m.: Team red’s leaving from Pooh’s corner

2:35 p.m.: Team red’s arrival at the Hab

2:40 p.m.: Beginning of the emergency procedures

3:40 p.m.: End of the emergency procedures

3:45 p.m.: EVA#7 back in the Hab

EVA Report – February 26th

Dear CapCom,

Here is the EVA Report:

26 Feb 2018

EVA #6 Crew members: Jérémy Auclair (EVA leader), Victoria Da-Poïan, Gabriel Payen, Alexandre Martin


– Near Pooh’s Corner: 12S-518900E-4252000N

– White Rock Canyon: 12S-520200E-4247500N

Time: departure at 09:09 a.m. Duration: 2 hour 13 minutes

Narrative: My first EVA as EVA leader went really well ! We changer the battery that powered the LOAC instrument (the old one was completely empty), we tested the battery powering the MegaARES instrument and confirmed it needed to be changed quickly (tomorrow). Otherwise both instruments are supporting the desert conditions perfectly. We then went to white rock canyon for exploration and photography. The walk went perfectly, the temperature was perfect, our reparations on the back packs prevented the apparition of fog on our helmets. We started to walk inside the canyon and on our way back we walk on the ridge (at a safe distance) to change the point of view. We finished the walk around 11am. Having nothing else to do we went back to the Hab to refill the ATVs and end the EVA. Everything went fine.


– LOAC: Maintenance – replacement of the battery

– MegaARES: Maintenance – Battery level test

– White Rock Canyon – Exploration and photography

Vehicles: 1 Rover (Curiosity), 2 ATVs (350.1, 350.2)

9:09 a.m.: EVA#6 left Airlock

9:15 a.m.: EVA left the Hab using Curiosity, ATVs 350.1 and 350.2

9:23 a.m.: Arrival on the working place

9:25 a.m.: Beginning of LOAC battery replacement and MegaARES battery test

9:35 a.m.: End of maintenance

9:40 a.m.: Beginning of the drive for White Rock Canyon

10:00 a.m.: Arrival at White Rock Canyon and beginning of walk

10:50 a.m.: End of walk and beginning of the drive back to the Hab

11:05 a.m.: Arrival at the Hab

11:15 a.m.: Entry in the Airlock



Commander of Crew 189

EVA Report

Hello again CapCom,

Here is today’s EVA report:

EVA Report:
EVA #3 Crew members: Louis Mangin (EVA leader), Benoit Floquet, Laurent Bizien, Alexandre Martin

Author: Louis Mangin
Location: Around the Hab 12S-518100E-4250700N
Near Pooh’s Corner: 12S-518900E-4250900N
Cactus road: 12S-52200E-4252700N

Time: departure at 9:30 p.m. Duration: 2 hours

– Deploy the LOAC after reparation
– Check MegaARES status
– Explore around Cactus Road
– Shoot Media photos

We started by getting a lot of gear out of the engineering airlock: wastes, LOAC and tools. We then struggled to start the ATVs because of the coldness. When arrived on site, I had fog on my helmet because of the walk carrying the device, so that I didn’t see exactly chat happened, but while plugging the power wire, the power system overheated, and wires melted. I then disconnected the battery, but it was already too late. We evaluated then that only the power system was touched and apparently were right to guess so. Anyway, we packed everything back and went back to the hab to bring the sensor back. At the same time Benoit had checked the MegaARES and everything was fine there. After letting the broken LOAC in the engineering airlock, we headed to Cactus Road. We then climbed a hill, went to an entrance of Candor Chasma and shoot some group photos with our school’s banner. The fog intensified for the ones wearing the first version of the suits, and it was pretty cold, so that we headed back to the Hab early.

9:30 a.m.: EVA#3 left Airlock
9:40 a.m.: EVA left the Hab using two ATVs and one rover
9:45 a.m.: Arrival at place of interest
10:00 a.m.: Failure due to LOAC’s power system overheating
10:20 a.m.: Back to Hab to put back the LOAC
10:30 a.m.: Leaving the Hab to Cactus Road
10:40 a.m.: Exploration of Cactus Road surroundings
11:15 a.m.: Leaving Cactus Road
11:30 a.m.: EVA End

Louis Mangin

EVA Report – February 20th

EVA Report:

EVA #2 Crew members: Victoria Da-Poian (EVA leader), Jérémy Auclair, Gabriel Payen, Benoit Floquet

Location: – Around the Hab 12S-518100E-4250700N

Time: departure at 09:30 a.m. Duration: 2 hour 40 minutes

Narration: I was really happy to explore again the red cliffs with my new crew. Since this morning, everyone has experienced the EVAs. We managed to deploy one experiment (the MegaARES) but a small problem interrupted the LOAC deployment. We will fix it with Jérémy in order to try again tomorrow. Even if Gabriel had a very foggy helmet (fans problems), and everyone seems really happy and I have to say the photos Benoit took are really nice. During our EVA, we had a very nice teamwork. I am very proud of our common work and I am looking forward to going back to EVA with them.

EVA Report – February 19th

Hello Graeme,

Here is our first EVA’s reports. We experienced a better weather than expected so that we were able to stay longer on the outside and visit the surroundings!

EVA Report:
EVA #1 Crew members: Louis Mangin (EVA leader), Jérémy Auclair, Alexandre Martin,

Location: Around the Hab 12S-518100E-4250700N
Near Pooh’s Corner: 12S-518900E-4250900N
Under Eastern Kissing Camel Ridge: 12S-518500E-4249700N
Top of Hab Ridge: 12S-517800E-4249700N
Zubrin’s head: 12S-519500E-4248100N

Time: departure at 2:00 p.m. Duration: 1 hour 50 minutes

– Find a location to deploy our two atmosphere study devices – First EVA logger in situ test
– Exploration and EVA training for newcomers
– First photos and videos shots

1:58 p.m.: EVA#1 left Airlock
2:01 p.m.: EVA left the Hab using Deimos & Opportunity
2:09 p.m.: Arrival on first place of interest: Pooh’s Corner 2:40 p.m.: Leaving Pooh’s Corner
2:50 p.m.: Arrival on second place of interest
3:00 p.m.: Climbing to the top of Hab Ridge (third place of interest) 3:25 p.m.: Getting back on the rovers
3:35 p.m.: Going to Zubrin’s head to shoot photos
3:50 p.m.: Getting back in the Hab


Louis Mangin
Crew 189 Commander

EVA Report – February 9th

Crew 188 EVA #14 Summary Report 09FEB2018

Author’s name: Julia DeMarines, Crew Astronomer, Green Hab Operator

Purpose of EVA: To perform additional video filming using either the GoPRO or Insta360 PRO camera on the ridge around the Hab. While walking, we will collect rock and plant samples for in-situ chlorophyll samples for analysis.

Location of EVA: Hab Ridge
UDM27 Coordinates: 12S, 5175750E, 54250500N

Number of EVA Crew: 3

Participants: Ryan Kobrick, Sarah Jane Pell, Julia DeMarines

EVA Commander: Julia DeMarines

Road(s) and routes per MDRS map:

1) Exit Hab and head to the hills directly west to the Hab

2) Walk to the Sagan Road

3) Climb to summit Hab Ridge

4) Follow the road on Hab Ridge Road as time allows

Mode of Travel: Walking

Vehicles you will be using (if applicable): None.

Duration: 2 hours 41 minutes

EVA Departure Time: 10:26

EVA Return Time: 13:07


We left the hab a little after the predicted time due to two last minute bathroom needs. The weather was sunny with some haze and no wind. The crew soon felt overheated, helmets/glasses were fogging up and had to stop frequently to let the sun burn off the fog. Once on the ridge, Julia selected two sites for chlorophyll detection samples, and Sarah Jane set up her Insta360º Pro camera for three different shots. Ryan recorded the journey on a GoPro camera to document the experience. Once our objectives were accomplished we turned around and walked back to the hab.

EVA Report – February 9th

EVA #15 EVA Report 09Feb2018

Author’s name: Tatsunari Tomiyama AHFP, Health & Safety Officer

Purpose of EVA: To collect micrometeorite soil samples at the Tooth Ridge and take photo and video for the mission documentary.

Location of EVA: Toothy Ridge

UDM27 Coordinates: 12 S, 516871 E, 4255884 N

Number of EVA Crew: 3

Participants: Zac Trolley, Tatsunari Tomiyama, Renee Garifi

EVA Commander: Renee Garifi

Rovers used: ATV

1, 2, 3

Duration: 1hours 29minutes

EVA Departure Time: 14:16

EVA Return Time: 15:45


The objective of this EVA was to collect meteorite samples at the toothy ridge. We prepared and entered the airlock in timely manner. However, all crew members have issues breathing while walking. Therefore, we decided to return to avoid a critical incident. The details on our breathing issues for crew members will be reported in the operation report. Please reference the operation issue report as a detailed summary.